A 206-kg crate is pushed horizontally with a force of 705 N. If the coefficient of friction is 0.20, calculate the acceleration of the crate.
1
Expert's answer
2012-10-24T10:34:44-0400
The friction force acting on a crate is

F(friction) = μ*m*g = 0.20*206[kg]*9.81[m/s²] = 404.17 [N].

The net force acting on a crate is

F(net) = F(pushing) - F(friction) = 705[N] - 404.17[N] = 300.83 [N].

According to the second Newton's law the acceleration of a crate is

a = F(net)/m = 300.83[N]/206[kg] = 1.46 [m/s²].